### v1

# Binaural Target Speaker Extraction using HRTFs and a Complex-Valued Neural Network

Abstract
In this work, we aim to imitate the human ability to selectively attend to a single speaker, even in the presence of multiple simultaneous talkers. To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker. Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets in different languages. We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals. This deviates from conventional approaches that use spectrograms or treat the real and imaginary components of the STFT as separate real-valued inputs. We begin by evaluating the method in an anechoic, noise-free scenario, where it demonstrates excellent extraction performance while effectively preserving the binaural cues of the target signal. Next, we test a modified variant under mild reverberation conditions. This version remains robust in reverberant conditions, maintaining speech clarity, preserving source directionality, and simultaneously reducing reverberation. Demo-page: https://bi-ctse-hrtf.github.io
Index Terms:
Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
I Introduction
Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking. Such a scenario is often referred to as the cocktail party problem [1]. While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments. Binaural Target Speaker Extraction (Bi-TSE) aims at selectively extracting one desired speaker of the multiple competing speakers, in a binaural setting. In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3]. Most hearing aids feature multiple microphones on each earpiece and fall into one of the two categories: 1) Bilateral systems, in which each device operates independently, yielding a single-channel output. 2) Shared systems, where both earpieces share all microphones, producing a dual-channel output.
Two important spatial cues are the Interaural Level Difference (ILD) and Interaural Time Difference (ITD), which capture magnitude and arrival-time differences between the left and right ears. In bilateral systems, where each device processes its own channel independently, it is particularly challenging to preserve these binaural cues. In contrast, Multi input Multi output (MIMO) solutions jointly process both channels, yielding more natural and spatially accurate outputs. For example, [4] presents a Binaural Linearly Constrained Minimum Variance (BLCMV) beamformer that separates sources while maintaining binaural cues. More recent work, such as [5], employs a Deep Neural Network (DNN) to process all microphone signals and produce left- and right-channel outputs, effectively preserving spatial information.
Unlike Blind Source Separation (BSS) algorithms, which separate all sources, Target Speaker Extraction (TSE) algorithms leverage prior knowledge of the target speaker(s) to isolate their voice. The prior knowledge about the speaker—commonly referred to as “clue” [6]—can include any additional information about the target speaker. For instance, it might be visual data depicting the speaker’s appearance as in [7] that uses a face image to extract its corrospondig speech, or [8] that uses both audio and visual clues, where the audio clue is an enrollment utterance and the visual clue is a video of the speaker speaking within the mixture. Another popular clue is a pre-recorded sample used to extract speaker-specific features, as in [9, 10, 11, 12].
While examining prior works, such as [13], we observed the particular challenge of separating two voices with highly similar characteristics—for example, speakers of the same gender. Consequently, we opted for an approach that does not depend on speaker–embedding–based clues. Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better. For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA. These spatially guided methods consistently outperform single-microphone solutions.
In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that leverages the listener’s HRTF, aligned with the target speaker’s DOA, as an extraction clue. Our approach employs a complex‑valued DNN that operates directly in the STFT domain. Unlike real‑valued models that treat real and imaginary parts separately—or architectures that rely solely on the magnitude spectrogram—our network processes the full complex representation, more effectively preserving phase information. The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved. Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics. The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases.
II Problem Formulation
We address the problem of multiple concurrent speakers, captured by two microphones, each mounted on one of the listener’s ears in a noiseless environment. For clarity, we focus on the simpler case of two concurrent speakers, as formulated below.
Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation). Similarly, the interfering speaker’s DOA is ($\theta_{i}$,$\phi_{i}$). Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively. This signal can be written as
$\bm{x}_{b}(n,k)\;=\;s_{d}(n,k)\,\bm{h}(\theta_{d},\phi_{d},k)+s_{i}(n,k)\,\bm{h}(\theta_{i},\phi_{i},k)$ (1)
where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech of the desired speaker and the interfering speaker, respectively. Let $\gamma=\{d,i\}$ denote the speaker index, either the desired ($d$) or the interferer ($i$). Here, $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ is a Transfer Function (TrF) modeling acoustic propagation from source $\gamma$ to both earpieces as a superposition of HRTFs associated with the DOAs of all reflections. The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, represents the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$. According to the precedence effect [16], the direct-path HRTF dictates the perceived source DOA. Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity. Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$ Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
Our work has three main objectives: 1) to extract the target speaker at a known DOA from binaural input using the listener’s HRTF; 2) to generate a binaural output that preserves the spatial cues of the desired source; and 3) to dereverberate the output, retaining only the direct-path HRTF.
III Proposed Method
We propose a deep neural network model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task. Given the HRTF from the desired source DOA $(\theta_{d},\phi_{d})$, namely $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$. We denote the extracted signal as $\hat{\tilde{\bm{s}}}_{d}$.
III-A Architecture
The proposed model architecture is based on the approach presented in [14], modified to fully handle complex-valued inputs, rather than the real and imaginary components independently.
Encoder-Decoder Blocks of the U-net: The encoder comprises five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation function:
$\mathbb{C}\text{ReLU}(z)=\text{ReLU}\left(\mathcal{R}(z)\right)+\mathrm{j}\cdot\text{ReLU}\left(\mathcal{I}(z)\right).$ (2)
Here $\text{ReLU}(x)=\max(0,x)$ and $\mathcal{R}(z),\mathcal{I}(z)$ denote the real and imaginary parts of z, respectively. Additionally, $\mathrm{j}$ represents the imaginary unit. Similarly, the decoder comprises five up-convolution blocks, each employing transposed convolutional layers. Skip connections are utilized between corresponding encoder and decoder blocks of the U-net structure to improve reconstruction accuracy. Additionally, the decoder includes a final convolutional layer with no activation function, enabling the representation of both positive and negative values at the output. This design choice is essential since the STFT representation inherently includes both positive and negative real and imaginary components.
Bottelneck Processing: The extraction clue is applied to the encoded input features at the bottleneck, with self-attention layers employed both before and after its application. However, unlike traditional methods, we implement these self-attention layers directly on complex-valued tensors by replacing the softmax activation function with the complex-valued $\mathbb{C}\text{ReLU}$.
To capture cross-channel (left and right) information in the HRTF, we apply the self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage. A block diagram of this architecture is presented in Fig. 1.
[…]
Figure 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed signal at the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of estimated desired signal.
III-B Loss function
We initially trained the proposed model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
[…]
where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all samples of the target and estimated signals, respectively. However, relying solely on the SI-SDR loss proved challenging in preserving accurate spectral details, resulting in noticeable musical noise artifacts in the extracted signals. Musical noise refers to tonal distortions that occur due to inaccuracies or rapid variations in the estimated signal’s frequency-domain representation [17]. Such artifacts are commonly observed in signal enhancement tasks when spectral continuity and consistency are not adequately enforced. To mitigate these artifacts, we introduced an additional frequency-domain loss, the $L1$- Mean Absolute Error (MAE) loss:
$\mathcal{L}_{\text{MAE}}=\frac{1}{2KN}\sum_{c=\{l,r\}}\sum_{k=0}^{K-1}\sum_{n=0}^{N-1}\left||\tilde{s}_{d}^{c}(n,k)|-|\hat{\tilde{s}}_{d}^{c}(n,k)|\right|.$ (4)
Here, $c=\{l,r\}$ denotes the channels, left or right. To improve data utilization, both speakers were extracted from each mixed signal in each training step by alternating the HRTF clues and averaging the SI-SDR losses. The overall training loss is therefore defined in (5), as a weighted combination of the SI-SDR and $L1$-MAE losses:
$\mathcal{L}=\frac{1}{2}\sum_{\gamma=\{d,i\}}\left(-\text{SI-SDR}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})+\alpha\cdot\mathcal{L}_{\text{MAE}}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})\right).$ (5)
Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a two-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$). Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality.
IV Experimental Study
In this section, we describe the data-generation process and experimental setup, evaluate the performance of the proposed approach, and compare it with previous TSE methods.
IV-A Data Generation
Binaural Impulse Rsponses: The binaural Impulse Responses are formed by superimposing attenuated and time-shifted HRIRs, corresponding to acoustic reflections generated via the image method [18], as expressed in (6).
$\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)=\sum_{\ell=0}^{L}a_{\ell}\cdot\bar{\bm{h}}_{\text{hrir}}(\theta_{\ell},\phi_{\ell},t-\tau_{\ell})$ (6)
where $L$ is the reflection order. The direct path corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$.
The RIEC HRTF Dataset [19, 20]: A measured HRTF database consists of 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA.
-90
-80
-70
-60
-50
-40
-30
-20
-10
0
10
20
30
40
50
60
70
80
90
Refer to captionRefer to captionRefer to caption
Figure 2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$ [^2] Images sourced from: https://www.freepik.com
We employed the SofaMyRoom framework [23] to simulate reverberant and anechoic conditions, similar to (6). This tool places a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a target reverberation time. It also supports zero-order simulations for anechoic scenarios. For training, we sampled $T_{60}\sim\mathcal{U}[0.2,0.8]\,$s, and for testing, $T_{60}\sim\mathcal{U}[0.1,0.4]\,$s. Speaker and listener positions were also drawn uniformly within the room’s dimensions. The simulation arrangement is illustrated in Fig. 2.
Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [24]. To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
IV-B Algorithm Setup
The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, and transformed using a 512-sample STFT with 75% overlap. Our model was trained in a supervised manner. For the anechoic setting, the target signals were generated by convolving the desired speech signal with the corresponding HRTF. For the reverberant setting, the targets were obtained by simulating a reverberant room and using only the first arrival (i.e., zero-order reflection). The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and transformed to the frequency domain using Fast Fourier Transform (FFT).
IV-C Metrics
To evaluate our method’s performance, we used SI-SDR[27], Perceptual Evaluation of Speech Quality (PESQ)[28], and Short-time Objective Intelligibility (STOI)[29] as measures of speech quality and intelligibility. lFurthermore, to verify that the estimated signals retain their binaural characteristics, we compare the Probability Density Function (p.d.f.) of the ILD and ITD of the target signals and the extracted signals. The similarity of the p.d.f.s indicates that the spatial cues are preserved.
IV-D Competing Methods
To ensure a fair comparison, all baseline methods operate under an identical setup, namely, binaural TSE using only two microphones (one per ear). Accordingly, we compare our method against the BLCMV beamformer [4] and a recent DNN-based model [30], both designed specifically for this setup. In addition, we trained two variants of our model. The first is Bi-TSE-HRTF-RI, a real-valued model, in which the complex-value input features are split into Real-Imaginary (RI) components and concatenated, resulting in four input channels instead of the two input channels in the original scheme. The same procedure is applied to the HRTF clue. The second is Bi-TSE-DOA, a complex-valued model that incorporates the DOA as a clue instead of the HRTF. Here, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs and passed through a fully-connected layer to match the dimensionality of the bottleneck. For the specified variants, both terms of (5) remain intact, i.e., the target signal is the binaural rendering of the desired speech signal with the correct HRTF.
(a) SI-SDR[dB]
(b) PESQ
(c) STOI
Figure 3: Histograms of speech quality and intelligibility measures, before and after the application of the proposed TSE algorithm for the anechoic setting.
TABLE I: The evaluation datasets include 3k speech samples mixed with SIR sampled from $\mathcal{U}[0,5]$ dB. Speaker overlap is either sampled from $\mathcal{U}[0\%,100\%]$ (designated by) or set to 100% (designated by∗∗), with audio lengths of 5 s and 8 s, respectively. Best scores per metric and dataset are highlighted in bold.
Method Dataset Reverberant SI-SDR (dB) $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ $\Delta$ILD [dB] $\downarrow$ $\Delta$ITD [mS] $\downarrow$
Unprocessed LS∗∗ 2.4 1.36 0.78 - -
BLCMV Beamformer [4] LS∗∗ 0.86 3.93 0.95 0.7341 0.6032
Bi-CSim-TSE [30] LS 18.52 3.05 0.96 - -
Bi-TSE-HRTF (proposed) LS∗∗ 19.35 3.27 0.98 0.1311 0.0033
Unprocessed MLS-French∗∗ 2.37 1.29 0.77 - -
BLCMV Beamformer [4] MLS-French∗∗ 0.98 3.92 0.95 0.7092 0.5829
Bi-TSE-HRTF (proposed) MLS-French∗∗ 18.9 3.23 0.98 0.2697 0.0120
Unprocessed WSJ0∗∗ 0.06 1.23 0.73 - -
BLCMV Beamformer [4] WSJ0∗∗ 1.51 3.98 0.96 0.733 0.5669
Bi-TSE-DOA WSJ0∗∗ 16.3 2.47 0.97 0.1304 1.0145
Bi-TSE-HRTF-RI WSJ0∗∗ 21.36 2.97 0.99 0.0934 0.00435
Bi-TSE-HRTF (proposed) WSJ0∗∗ 20.37 3.49 0.99 0.1003 0.0027
Unprocessed WSJ0∗∗ -1.15 1.21 0.7 - -
BLCMV Beamformer [4] WSJ0∗∗ 1.37 1.6 0.82 0.3303 0.5439
Bi-TSE-DOA WSJ0∗∗ 9.53 2.08 0.93 0.2410 1.0125
Bi-TSE-HRTF (proposed) WSJ0∗∗ 10.37 2.60 0.95 0.1998 0.0036
IV-E Results
In Fig. 3, we show histograms of the SI-SDR, PESQ, and STOI metrics before and after applying our TSE algorithm in the anechoic scenario. All three distributions shift markedly to the right, demonstrating substantial gains across signal quality, perceptual quality, and intelligibility. This confirms the algorithm’s effectiveness in isolating the desired speaker.
In Table I, we compare our method against competing algorithms across all datasets under both anechoic and reverberant conditions. Our approach achieves significant gains on every metric except PESQ, where the BLCMV beamformer slightly outperforms the proposed method in anechoic case. This may be attributed to its distortionless response design criterion.
When comparing the two additional variants of our model, denoted as Bi-TSE-DOA and Bi-TSE-HRTF-RI in Table I, we observe that, in contrast to the HRTF-guided model, the DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and is unable to preserve the binaural cues. We evaluated this model under reverberant conditions to emphasize that the HRTF is intrinsically tied to our problem formulation. The superior performance of the HRTF-guided network supports these findings. On the other end, the RI model achieves comparable SI-SDR and binaural cue preservation to the complex-valued model but suffers from lower output quality, as shown by its lower PESQ score in the anechoic setting.
Remarkably, our approach demonstrates cross-dataset generalizability: although our model was trained exclusively on WSJ0-CSR1, it performs exceptionally well on other corpora, including a French dataset. We attribute this generalization to the absence of speaker-specific priors in our approach.
Figure 4: A polar plot showing the SI-SDR of the output signals for each DOA, with the desired speaker located at $\theta_{d}=-75^{\circ}$ (red) and the interfering speaker at $\theta_{i}=-10^{\circ}$ (blue). The bold circle represents the 0 dB SI-SDR threshold.
When inspecting the extracted signals, we found that the model effectively enhances the target speaker while nearly eliminating the interferer. To quantify its spatial resolution, we devised a “scanning” procedure: we iteratively extract signals using HRTFs corresponding to azimuth angles in $[-90^{\circ},90^{\circ}]$ at a fixed elevation angle and compute the SI-SDR of each extraction against the true target. Plotting these values yields a beam-pattern-like diagram (Fig. 4), which shows a pronounced peak at the correct DOA. Additionally, the width of the main lobe indicates that the model remains effective even when the HRTF clue deviates slightly from the true HRTF. The proposed method’s ability to preserve binaural cues is also evaluated and compared with the BLCMV, where applicable. Table I presents the mean ILD and ITD deviations. Figure 5 depicts the p.d.f.s of both cues for the target and extracted signals, showing a close match and confirming faithful spatial cue preservation.
(a) Mixture
(b) $\hat{\tilde{\bm{s}}}_{d}$
(c) $\hat{\tilde{\bm{s}}}_{i}$
Figure 5: The p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=90^{\circ}$, $\theta_{i}=15^{\circ}$, and $\phi_{d,i}=0^{\circ}$. Graphs produced by [31].
V Conclusions
This paper presents a speaker‑independent TSE method that leverages the listener’s HRTFs in both anechoic and reverberant settings, resulting in a personalized solution for the proposed problem. Through cross‑dataset evaluations, we demonstrate its robust generalization across diverse corpora. Our approach not only delivers high‑quality speech but also faithfully preserves the binaural character of the signal. Comparisons with other binaural TSE techniques underscore the importance of spatial cues tied to the listener’s HRTF for accurate and consistent target‑speaker extraction under varying acoustic conditions.