### v2

# Binaural Target Speaker Extraction using HRTFs

Abstract
In this work, we aim to imitate the human ability to selectively attend to a single speaker, even in the presence of multiple simultaneous talkers. To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker. Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets and languages. We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals, and compare it to a Real-Imaginary (RI)-based neural network, demonstrating the advantages of the former. We first evaluate the method in an anechoic, noise-free scenario, achieving excellent extraction performance while preserving the binaural cues of the target signal. We then extend the evaluation to reverberant conditions. Our method proves robust, maintaining speech clarity and source directionality while simultaneously reducing reverberation. A comparative analysis with existing binaural Target Speaker Extraction (TSE) methods demonstrates that our approach attains performance on par with competing techniques in terms of noise reduction and perceptual quality, while offering a clear advantage in preserving binaural cues. Demo-page: https://bi-ctse-hrtf.github.io
Index Terms:
Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
I Introduction
Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking. Such a scenario is often referred to as the cocktail party problem [1]. While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments. Binaural Target Speaker Extraction (Bi-TSE) aims to selectively extract the desired speaker from multiple competing speakers in a binaural setting. In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3]. Most hearing aids employ either a single microphone or multiple microphones per earpiece and can be broadly categorized into two groups: 1) Bilateral systems, where each device operates independently and provides a single-channel output per earpiece; 2) Shared systems, where all microphones across both earpieces are jointly utilized, yielding a dual-channel output.
Two important spatial cues are the Interaural Level Difference (ILD) and Interaural Time Difference (ITD), which capture magnitude and arrival-time differences between the left and right ears. In bilateral systems, where each device processes its own channel independently, it is particularly challenging to preserve these binaural cues. In contrast, Multi input Multi output (MIMO) solutions jointly process both channels, yielding more natural and spatially accurate outputs. For example, [4] presents a Binaural Linearly Constrained Minimum Variance (BLCMV) beamformer that separates sources while maintaining binaural cues. More recent work, such as [5], employs a Deep Neural Network (DNN) to process all microphone signals and produce left- and right-channel outputs, effectively preserving spatial information.
Unlike Blind Source Separation (BSS) algorithms, which separate all sources, TSE algorithms leverage prior knowledge of the target speaker(s) to isolate their voice. This prior knowledge—commonly referred to as a “clue” [6]—can take various forms. A widely used clue is a pre-recorded speech sample, often referred to as an enrollment utterance, which provides speaker-specific features [7, 8, 9, 10]. Clues may also come from other modalities, such as visual data. For example, [11] uses a face image to extract the corresponding speech, while [12] combines audio and visual clues, with the audio clue being an enrollment utterance and the visual clue a video of the speaker speaking within the mixture.
While examining prior works, such as [13], we observed a particular challenge in separating two voices with highly similar characteristics, for example, speakers of the same gender. Consequently, we opted for an approach that does not depend on speaker–embedding–based clues. Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better. For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA and extracts spatial features from it. These spatially guided methods consistently outperform single-microphone solutions.
In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that exploits the listener’s HRTF, aligned with the target speaker’s DOA, as the primary extraction cue. Our approach employs a complex‑valued DNN that operates directly in the STFT domain. The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved. Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics. The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases, and compared with two competing methods, a beamformer-based solution and a DNN-based solution.
II Problem Formulation
We address the problem of multiple concurrent speakers, captured by two microphones, each mounted on one of the listener’s ears in a noiseless environment. For clarity, we focus on the simpler case of two concurrent speakers, as formulated below.
Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation). Similarly, the interfering speaker’s DOA is ($\theta_{i}$$\phi_{i}$). Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively. The binaural signal can be written as
$\bm{x}_{b}(n,k)\;=\;s_{d}(n,k)\,\bm{h}(\theta_{d},\phi_{d},k)+s_{i}(n,k)\,\bm{h}(\theta_{i},\phi_{i},k)$ (1)
where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech signals of the desired speaker and the interfering speaker, respectively. Let $\gamma\in\{d,i\}$ denote the speaker index, where $d$ refers to the desired source and $i$ to the interferer. The acoustic transfer from source $\gamma$ to both earpieces is described by $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, which models the Transfer Function (TrF) as a superposition of HRTFs corresponding to the DOAs of all reflections. The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, characterizes the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$.
According to the precedence effect [16], the direct-path HRTF dictates the perceived source DOA. Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity. Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$. Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
Given the HRTF corresponding to the desired source’s DOA $(\theta_{d},\phi_{d})$, namely the binaural response $\bm{h}_{\text{hrtf}}(\theta_{d},\phi_{d},k)$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$. The extracted signal is denoted by $\hat{\tilde{\bm{s}}}_{d}$. This binaural output is required to: 1) preserve the spatial cues of the desired source, and 2) dereverberate the output by retaining only the direct-path HRTF, thereby also dereverberating the desired speech signal.
III Proposed Method
We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task.
III-A Architecture
The proposed model architecture is based on the approach presented in [14], modified to fully handle dual-channel, complex-valued inputs, rather than the real and imaginary components independently. The proposed model employs a U-Net architecture, incorporating the HRTF at the bottleneck. The individual modules comprising the architecture are described in detail below.
Encoder-Decoder Blocks of the U-net: The encoder comprises five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation function:
$\mathbb{C}\text{ReLU}(z)=\text{ReLU}\left(\mathcal{R}(z)\right)+\mathrm{j}\cdot\text{ReLU}\left(\mathcal{I}(z)\right).$ (2)
Here $\text{ReLU}(x)=\max(0,x)$ and $\mathcal{R}(z),\mathcal{I}(z)$ denote the real and imaginary parts of z, respectively. Additionally, $\mathrm{j}$ represents the imaginary unit. Similarly, the decoder comprises five up-convolution blocks, each employing transposed convolutional layers. Skip connections are utilized between corresponding encoder and decoder blocks of the U-net structure to improve reconstruction accuracy. Additionally, the decoder includes a final convolutional layer with no activation function, enabling the representation of both positive and negative values at the output. This design choice is essential since the STFT representation inherently includes both positive and negative real and imaginary components.
Bottelneck Processing: The extraction clue is applied to the encoded input features at the bottleneck, with self-attention layers employed both before and after its application. However, unlike traditional methods, we implement these self-attention layers directly on complex-valued tensors by replacing the softmax activation function with the complex-valued $\mathbb{C}\text{ReLU}$.
To capture cross-channel (left and right) information in the HRTF, we apply a self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage. A block diagram of the proposed architecture is presented in Fig. 1.
[…]
Fig. 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated binaural desired signal.
III-B Loss function
We initially trained the proposed model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
$\text{SI-SDR}(\bm{s},\hat{\bm{s}})=20\log_{10}\left(\frac{\|\beta\bm{s}\|}{\|\beta\bm{s}-\hat{\bm{s}}\|}\right);\quad\beta=\frac{\hat{\bm{s}}^{\top}\bm{s}}{\|\bm{s}\|^{2}}\,$, (3)
where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all samples of the target and estimated signals, respectively. However, relying solely on the SI-SDR loss proved challenging in preserving accurate spectral details, resulting in noticeable musical noise artifacts in the extracted signals. Musical noise refers to tonal distortions that occur due to inaccuracies or rapid variations in the estimated signal’s frequency-domain representation [17]. Such artifacts are commonly observed in signal enhancement tasks when spectral continuity and consistency are not adequately enforced. To mitigate these artifacts, we introduced an additional frequency-domain loss, the $L1$- Mean Absolute Error (MAE) loss:
$\mathcal{L}_{\text{MAE}}=\frac{1}{2KN}\sum_{c=\{l,r\}}\sum_{k=0}^{K-1}\sum_{n=0}^{N-1}\left||\tilde{s}_{d}^{c}(n,k)|-|\hat{\tilde{s}}_{d}^{c}(n,k)|\right|.$ (4)
Here, $c=\{l,r\}$ denotes the channels, left or right. To improve data utilization, both speakers were extracted from each mixed signal in each training step by alternating the HRTF clues and averaging the SI-SDR losses. The overall training loss is therefore defined in (5), as a weighted combination of the SI-SDR and $L1$-MAE losses:
$\mathcal{L}=\frac{1}{2}\sum_{\gamma=\{d,i\}}\left(-\text{SI-SDR}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})+\alpha\cdot\mathcal{L}_{\text{MAE}}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})\right).$ (5)
Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a multi-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$). Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality. The final stage fine-tunes the model using only the SI-SDR loss, as the $L_{1}$-MAE loss is prone to overfitting to the training data.
IV Experimental Study
In this section, we describe the data-generation process and experimental setup, evaluate the performance of the proposed approach, and compare it with previous TSE methods.
IV-A Data Generation
Binaural Impulse Rsponses: The binaural Impulse Responses are formed by superimposing attenuated and time-shifted HRIRs, corresponding to acoustic reflections generated via the image method [18], as expressed in (6).
$\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)=\sum_{\ell=0}^{L}a_{\ell}\cdot\bar{\bm{h}}_{\text{hrir}}(\theta_{\ell},\phi_{\ell},t-\tau_{\ell})$ (6)
where $L$ is the reflection order. The direct path corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$.
The RIEC HRTF Dataset [19, 20]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA.
We employed the SofaMyRoom framework [23] to simulate both anechoic and reverberant conditions, following the formulation in (6). This framework positions a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a specified reverberation time. It also supports zero-order simulations for anechoic scenarios. The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and both speaker and listener positions were drawn uniformly within the room. The overall simulation setup is illustrated in Fig. 2.
Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [24]. To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
IV-B Algorithm Setup
The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, then transformed with a 512-sample STFT (75% overlap). The model is trained in a supervised manner: in the anechoic setting, targets are generated by convolving the desired speech with the corresponding HRTF; in the reverberant setting, by simulating a room and keeping only the first arrival (zero-order reflection). The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and converted to the frequency domain via Fast Fourier Transform (FFT).
-90
-80
-70
-60
-50
-40
-30
-20
-10
0
10
20
30
40
50
60
70
80
90
Refer to captionRefer to captionRefer to caption
Fig. 2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$. Images source: freepik.com
IV-C Metrics
To evaluate our method, we employed SI-SDR [27], wide-band and narrow-band Perceptual Evaluation of Speech Quality (PESQ[28], and Short-time Objective Intelligibility (STOI[29] as intrusive metrics of speech quality and intelligibility, i.e., measures that compare the extracted signal against a reference target. In addition, we used DNS-MOS [30], a non-intrusive DNN-based predictor of Mean Opinion Score (MOS), which estimates speech quality (SIG), background noise quality (BAK), and overall quality (OVRL).
[…]
(a) Reverberant Mixture
(b) $\tilde{\bm{s}}_{d}$
(c) $\hat{\tilde{\bm{s}}}_{d}$
(d) $\tilde{\bm{s}}_{i}$
(e) $\hat{\tilde{\bm{s}}}_{i}$
Ground-truthBi-TSE-HRTF (proposed)BDE-BiTSE[15]BLCMV Beamformer [4]
Fig. 3: The joint p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=-60^{\circ}$, $\theta_{i}=-90^{\circ}$, and $\phi_{d,i}=-10^{\circ}$, with $T_{60}=0.63$ s, for the frequency band centerd at 500 Hz, where the ILD is less pronounced. $\tilde{\bm{s}}_{d}$ and $\tilde{\bm{s}}_{i}$ denotes the anechoic target signals. Graphs produced by [31].
Table I: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples. Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap . Best scores are highlighted in bold.
[…]
LS Unprocessed (2.688\3.380\3.140) -0.01 1.23 1.61 0.72 - -
BLCMV Beamformer [4] (2.827\3.304\3.495) -1.0 3.46 3.9 0.95 0.672 0.599
BDE-BiTSE [15] (3.072\3.471\3.768) 24.01 3.66 4.02 0.99 0.142 0.485
Bi-TSE-HRTF (proposed) (2.999\3.398\3.742) 20.18 3.14 3.78 0.98 0.1024 0.0042
MLS-French Unprocessed (2.499\3.252\2.947) -0.01 1.23 1.63 0.71 - -
BLCMV Beamformer (2.761\3.277\3.384) -1.03 3.37 3.82 0.94 0.694 0.608
BDE-BiTSE (2.987\3.403\3.709) 23.67 3.67 4.04 0.99 0.3 0.503
Bi-TSE-HRTF (proposed) (2.900\3.324\3.655) 19.86 3.14 3.79 0.98 0.1039 0.0041
WSJ0 Unprocessed (2.845\3.376\3.411) 0 1.25 1.57 0.72 - -
BLCMV Beamformer (3.044\3.444\3.748) -0.42 3.48 3.84 0.95 0.745 0.613
BDE-BiTSE (3.230\3.543\3.951) 25.59 3.77 4.01 1.0 0.126 0.484
Bi-TSE-DOA (2.965\3.376\3.660) 16.44 2.43 3.09 0.97 0.165 0.990
Bi-TSE-HRTF-RI (3.110\3.504\3.764) 21.39 2.93 3.57 0.99 0.085 0.0044
Bi-TSE-HRTF (proposed) (3.18\3.501\3.911) 21.41 3.47 3.88 0.99 0.0035 0.0756
WSJ0 (Reverb) Unprocessed (2.348\3.069\2.783) -1.9 1.15 1.44 0.69 - -
BLCMV Beamformer (1.840\2.612\2.038) -1.11 1.35 1.88 0.74 0.430 0.593
BDE-BiTSE (3.141\3.438\3.952) 14.8 3.09 3.56 0.98 0.204 0.386
Bi-TSE-DOA (2.681\3.196\3.324) 8.29 1.82 2.44 0.91 0.296 0.990
Bi-TSE-HRTF-RI (3.030\3.362\3.838) 11.18 2.01 2.74 0.95 0.174 1.001
Bi-TSE-HRTF (proposed) (2.86\3.348\3.481) 10.67 2.17 2.79 0.94 0.186 0.0082
IV-D Competing Methods
To ensure a fair comparison, all baseline methods were evaluated under an identical setup, namely binaural TSE with only two microphones (one per ear). Accordingly, we compared our method against the BLCMV beamformer [4] and a recent DNN-based model [15], both specifically designed for this scenario. The competing DNN, referred to as BDE-BiTSE, was trained following the procedure described in the paper, using the same training data as ours. We ensured that the strongest variant reported in the paper (BDE+CDF+IPD+SDF) was used for comparison.
In addition, we trained two variants of our model. The first, Bi-TSE-HRTF-RI, is a real-valued model in which the complex-valued input features are split into their RI components and concatenated, resulting in four input channels instead of two. The same decomposition is applied to the HRTF clue. The second, Bi-TSE-DOA, is a complex-valued model that incorporates the DOA as a clue instead of the HRTF. In this case, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs, and passed through a fully connected layer to match the dimensionality of the bottleneck. For both variants, the loss function in (5) remains unchanged, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.
IV-E Results
In Table I, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions. In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE. Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
Remarkably, both DNN-based methods demonstrate strong cross-dataset generalizability. Although the models were trained exclusively on WSJ0-CSR1, they perform well on other corpora, including a French dataset. We attribute this generalization to the absence of speaker-specific priors and to the reliance on a known HRTF database.
[…]
BLCMV Beamformer 1.7 >90>90 >150>150
BDE-BiTSE 1.0667 30 90
Bi-TSE-HRTF 0.0 -60 0
Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI. The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues, especially under reverberant conditions, underscoring the central role of HRTFs in our formulation. By contrast, the RI model achieves SI-SDR and binaural cue preservation comparable to the complex-valued model in the anechoic case, but suffers from reduced perceptual quality (lower PESQ) and fails to maintain cue preservation under reverberation.
Figure 3 shows the joint ITDILD p.d.f.s across methods. The close alignment of our model’s distribution with that of the target underscores its effectiveness in maintaining spatial cues. In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations. The dereverberation effect is also apparent, as the joint ITD-ILD p.d.f. becomes more concentrated, indicating dominance of the direct path in the output of the proposed method. Overall, it is evident that the proposed approach achieves this more effectively than competing techniques. Table LABEL:tab:delta_thata further highlights the spatial consistency observed in Fig. 3. Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec. IV-C. The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
Fig. 4: A polar plot of SI-SDR versus DOA, with the target at $\theta_{d}=-75^{\circ}$ (red) and the interferer at $\theta_{i}=-10^{\circ}$ (dashed blue). Stars mark true DOAs; the dotted arc denotes the 0,dB threshold.
Finally, to quantify the spatial resolution of our model, we employ a “scanning” procedure: signals are iteratively extracted using HRTFs at azimuths within $90^{\circ},90^{\circ}]$ (with fixed elevation), and the SI-SDR of each extraction is computed against the target. Plotting these values produces a beam-pattern-like diagram (Fig. 4) that exhibits a clear peak at the true DOA. The width of the main lobe further characterizes the model’s robustness when the HRTF cue deviates slightly from the ground-truth.
V Conclusions
This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, yielding a personalized solution to the task at hand. Through cross-dataset evaluations, we demonstrate robust generalization across diverse corpora. The proposed approach not only produces high-quality speech but also faithfully preserves the binaural characteristics of the signal. Comparative analyses with existing binaural TSE methods reveal substantial improvements in spatial consistency, while preserving high speech quality. These findings highlight the critical role of the listener’s HRTF in anchoring spatial cues, enabling accurate and consistent target-speaker extraction under a wide range of acoustic conditions.